OpenStudy (anonymous):

ABC is isosceles with AB=AC=8 units and BC=6 units. D and E are midpoints of AB and BC respectively. what is the length of DE

OpenStudy (anonymous):



OpenStudy (anonymous):

wait...

OpenStudy (anonymous):

\[\frac{ AD }{ AB } = \frac{ AE }{ AC } = \frac{ DE }{ BC }\]

OpenStudy (anonymous):

\[\frac{ 4 }{ 8 } = \frac{ 4 }{ 8 } = \frac{ DE }{ 6 }\]

OpenStudy (anonymous):

Ok?

OpenStudy (anonymous):

now find "DE"...you can...

OpenStudy (anonymous):

npe

OpenStudy (anonymous):

did you understand that how I found it?

OpenStudy (anonymous):

npe

OpenStudy (anonymous):

\[\Delta ADE \ is \ similar \to \Delta ABC\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then did you get it?

OpenStudy (anonymous):

no not the answer

OpenStudy (anonymous):

\[\frac{ 4 }{ 8 } = \frac{ DE }{ 6 } \to 24=8DE \to DE=\frac{ 24 }{ 8 }\]

OpenStudy (anonymous):

what is the first step in constructing a line parrallel to AB passing through an external point M, using only a compass and straightedge

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."